If vectors A and B are orthogonal, what is the component of B along the direction of A? What is the component of A along the direction of B?


1
Expert's answer
2021-05-04T12:09:05-0400

The component of "\mathbf{B}" along the direction of "\mathbf{A}" is given by the scalar product:


"B_A =\dfrac{ \mathbf{A}\cdot \mathbf{B}}{|\mathbf{\mathbf{B}}|}"

where "|\mathbf{B}|" is the magnitude of the vector. Since for the orthogonal vectors the scalar product is equal to 0 ("\mathbf{A}\cdot \mathbf{B} = 0" ). Thus, the component of "\mathbf{B}" along the direction of "\mathbf{A}" is zero.


Similarly, the component of "\mathbf{A}" along the direction of "\mathbf{B}" is:


"A_B =\dfrac{ \mathbf{B}\cdot \mathbf{A}}{|\mathbf{\mathbf{A}}|}"

Since "\mathbf{B}\cdot \mathbf{A} = \mathbf{A}\cdot \mathbf{B} =0" (the scalar product is commutative), the component of "\mathbf{A}" along the direction of "\mathbf{B}" is zero as well.


Answer. 0 and 0.